OpenStudy (anonymous):

simplify completely x^2+12x+35/3x+15?

OpenStudy (shamil98):

Factor the numerator and denominator.

OpenStudy (shamil98):

Once you do so, you will be able to cancel terms.

OpenStudy (anonymous):

x+5/3

OpenStudy (anonymous):

am i right @shamil98

OpenStudy (shamil98):

No.

OpenStudy (shamil98):

Factor x^2 + 12x + 35, what do you get?

OpenStudy (shamil98):

Factor 3x+15 what do you get? ( take the gcf)

OpenStudy (anonymous):

12 right? @shamil98

OpenStudy (shamil98):

No..

OpenStudy (shamil98):

\[\huge \frac{ x^2 + 12x + 35 }{ 3x + 15 } = \frac{ (x+7)(x+5) }{ 3(x+5) }\] Now simplify it..

OpenStudy (anonymous):

x^2+21/3

OpenStudy (shamil98):

No.

OpenStudy (anonymous):

waaaaaa i keep doing it wrong

OpenStudy (shamil98):

\[\huge \frac{ (x+7) \cancel{(x+5)} }{ 3\cancel{(x+5)} } = \frac{ x+7 }{ 3 }\] That is the simplified form.

OpenStudy (anonymous):

thnk you